# Fusionista2.0: Efficiency Retrieval System for Large-Scale Datasets

### v1

@@ 1 Introduction
As noted in prior works [Manning2008IR], effective information retrieval not only improves user experience but also supports decision-making processes across a wide range of domains. Consequently, research efforts must focus on developing retrieval methodologies with superior efficacy or refining and optimizing existing approaches to enhance their scalability within this domain and foster seamless interaction between the retrieval system and the user. Addressing these points, Video Browser Showdown [schoeffmann2012video], is an established, biennial competition that serves as a robust benchmark for assessing state-of-the-art interactive video retrieval systems. The VBS operates within realistic constraints, encompassing a spectrum of complex ad-hoc search issues, specifically: (1) Known-Item Search (KIS), (2) Ad-hoc Video Search (AVS), and (3) Visual Question Answering (VQA).
For VBS 2026, the primary dataset is the V3C [rossetto2019v3c], comprising over 28,000 videos with thousands of hours of content. Additional domain-specific datasets, such as Marine Video [marinevideo] and LapGynLHE [lapgynlhe], are also included. These heterogeneous datasets pose unique challenges and provide a comprehensive benchmark for testing the robustness and adaptability of retrieval systems.
In this paper, we present Fusionista2.0, an enhanced information retrieval system that extends Fusionista [fusionista] and Fustar [fustar] with multiple modalities, including text-to-image, image-to-image, object-level, and OCR/ASR-based retrieval. The main improvements are:
@@ 2.1 Overview
Our system, Fusionista2.0, builds upon the design principles and techniques of its predecessors, Fusionista [fusionista] and Fustar [fustar]. It provides a comprehensive set of retrieval functionalities tailored for the diverse query types in VBS2025. While these functionalities proved effective in VBS2025, scaling the system to multi-terabyte datasets revealed several challenges such as inference time, UX flow for thousand images. In Fusionista2.0, we introduce six key improvements (Figure 1), including:
@@ 2.2 Data Preparation and Extraction
In our previous system, the preprocessing pipeline was designed to maximize the accuracy of keyframe selection by leveraging multiple state-of-the-art models. Our previous system employed a preprocessing pipeline that combined CLIP-B/32 embeddings [radford2021], TransNetV2 [souvcek2020] scene detection, and clustering to select representative keyframes [tan2021], which were then used for multi-function search.
@@ 2.3 Textual Search
In our previous system, text-to-image retrieval relied on a single CLIP-based model, which often required a trade-off between efficiency and accuracy. In Fusionista2.0, we enhance this module by employing two state-of-the-art CLIP variants [clipbench]: CLIP-Sig400M and CLIP-ViT-5B. Both models are widely recognized for their balanced performance in terms of inference speed and retrieval accuracy, making them well-suited for large-scale interactive search scenarios.
@@ 2.4 OCR and ASR
In Fusionista2.0, we replace the previous PaddleOCR pipeline with the Vintern-1B-v3.5 model  [vintern]. This model, fine-tuned from InternVL2.5-1B [intervl], leverages low-resource languages datasets such as Viet-ShareGPT-4o-Text-VQA [vintern] and over one million OCR samples, achieving strong performance in many languages. Vintern-1B-v3.5 not only delivers high accuracy in scene text detection and recognition but also benefits from enhanced reasoning and general knowledge, enabling it to infer characters that are blurred or occluded.
While Whisper [whisper] provides excellent accuracy, its large model size makes it impractical for large-scale processing. Moreover, the majority of audio tracks in the VBS datasets contain ambient sounds rather than speech, reducing the necessity for high-capacity ASR models. For these reasons, we replace Whisper with faster_whisper [faster-whisper], a lightweight yet efficient variant optimized for speed. This choice significantly accelerates the transcription pipeline by four times [faster-whisper], while still maintaining sufficient accuracy to extract meaningful speech content.
@@ 2.5 Question Answering
InternVL-1B-Seq [intervl] 0.72 0.68 0.67 2.52
InternVL-1B-ffn6 [intervl] 0.79 0.74 0.71 2.48
InternVL-1B-ffn6-Seq [intervl] 0.84 0.79 0.75 4.80
LLaVA-0.5B-ffn6 [llava] 0.78 0.63 0.73 10.50
SmolVLM-0.5B-ffn6 [smovlm] 0.83 0.52 0.65 1.07
@@
Among all candidates, the InternVL-1B-ffn6-Seq [intervl] model demonstrated the most consistent performance across the three categories. It achieved an average inference time of less than 5 seconds per query over the entire dataset, while maintaining high accuracy in counting and frame-level information extraction tasks. This balance of speed and reliability makes it the most suitable choice for our system, ensuring accurate answers within the strict time limits of VBS.
@@ 2.6 Reranking
Our reranking approach is inspired by the methodology outlined in [aaaivqa4cir, summary], which utilizes a post-processing strategy that can be effortlessly incorporated into existing image retrieval systems. However, we extend this concept by designing a fully integrated AI-based VQA pipeline. In this pipeline, we employ GPT-4o [openai2024gpt4ocard] to formulate three yes-no questions based on the initial user query. These questions are crafted to explore the details of objects mentioned in the query, such as "Is there a dog in the scene?" or "Is the dog colored yellow?" Subsequently, both the questions and the associated images are evaluated by a vision language model to generate accurate responses. To enhance flexibility, we test multiple vision language models, including VideoLLaMA [DBLP:conf/emnlp/ZhangLB23_video_llama] and BLIP-2 [10.5555/3618408.3619222], leveraging the VLLMs[kwon2023efficient_vllm_lib] library to enable API-driven interactions.
@@ 3 UI/UX Upgrade
Our main key focus is UI/UX optimization, as efficient interaction directly impacts retrieval speed and usability. We redesigned the interface for responsiveness, accessibility (WCAG [WCAG21] compliance with shadcn/ui[shadcn], better feedback via loading states and error handling. Migration from Create React App (CRA) [cra] to Vite [vite] enhance significantly system performance.

### v2

@@ 1 Introduction
As noted in prior works [12], effective information retrieval not only improves user experience but also supports decision-making processes across a wide range of domains. Consequently, research efforts must focus on developing retrieval methodologies with superior efficacy or refining and optimizing existing approaches to enhance their scalability within this domain and foster seamless interaction between the retrieval system and the user. Addressing these points, Video Browser Showdown [19], is an established, biennial competition that serves as a robust benchmark for assessing state-of-the-art interactive video retrieval systems. The VBS operates within realistic constraints, encompassing a spectrum of complex ad-hoc search issues, specifically: (1) Known-Item Search (KIS), (2) Ad-hoc Video Search (AVS), and (3) Visual Question Answering (VQA).
For VBS 2026, the primary dataset is the V3C [18], comprising over 28,000 videos with thousands of hours of content. Additional domain-specific datasets, such as Marine Video [24] and LapGynLHE [14], are also included. These heterogeneous datasets pose unique challenges and provide a comprehensive benchmark for testing the robustness and adaptability of retrieval systems.
In this paper, we present Fusionista2.0, an enhanced information retrieval system that extends Fusionista [9] and Fustar [7] with multiple modalities, including text-to-image, image-to-image, object-level, and OCR/ASR-based retrieval. The main improvements are:
@@ 2.1 Overview
Our system, Fusionista2.0, builds upon the design principles and techniques of its predecessors, Fusionista [9] and Fustar [7]. It provides a comprehensive set of retrieval functionalities tailored for the diverse query types in VBS2025. While these functionalities proved effective in VBS2025, scaling the system to multi-terabyte datasets revealed several challenges such as inference time, UX flow for thousand images. In Fusionista2.0, we introduce six key improvements (Figure 1), including:
@@ 2.2 Data Preparation and Extraction
In our previous system, the preprocessing pipeline was designed to maximize the accuracy of keyframe selection by leveraging multiple state-of-the-art models. Our previous system employed a preprocessing pipeline that combined CLIP-B/32 embeddings [16], TransNetV2 [21] scene detection, and clustering to select representative keyframes [23], which were then used for multi-function search.
@@ 2.3 Textual Search
In our previous system, text-to-image retrieval relied on a single CLIP-based model, which often required a trade-off between efficiency and accuracy. In Fusionista2.0, we enhance this module by employing two state-of-the-art CLIP variants [2]: CLIP-Sig400M and CLIP-ViT-5B. Both models are widely recognized for their balanced performance in terms of inference speed and retrieval accuracy, making them well-suited for large-scale interactive search scenarios.
@@ 2.4 OCR and ASR
In Fusionista2.0, we replace the previous PaddleOCR pipeline with the Vintern-1B-v3.5 model  [3]. This model, fine-tuned from InternVL2.5-1B [1], leverages low-resource languages datasets such as Viet-ShareGPT-4o-Text-VQA [3] and over one million OCR samples, achieving strong performance in many languages. Vintern-1B-v3.5 not only delivers high accuracy in scene text detection and recognition but also benefits from enhanced reasoning and general knowledge, enabling it to infer characters that are blurred or occluded.
While Whisper [17] provides excellent accuracy, its large model size makes it impractical for large-scale processing. Moreover, the majority of audio tracks in the VBS datasets contain ambient sounds rather than speech, reducing the necessity for high-capacity ASR models. For these reasons, we replace Whisper with faster_whisper [22], a lightweight yet efficient variant optimized for speed. This choice significantly accelerates the transcription pipeline by four times [22], while still maintaining sufficient accuracy to extract meaningful speech content.
@@ 2.5 Question Answering
InternVL-1B-Seq [1] 0.72 0.68 0.67 2.52
InternVL-1B-ffn6 [1] 0.79 0.74 0.71 2.48
InternVL-1B-ffn6-Seq [1] 0.84 0.79 0.75 4.80
LLaVA-0.5B-ffn6 [10] 0.78 0.63 0.73 10.50
SmolVLM-0.5B-ffn6 [13] 0.83 0.52 0.65 1.07
@@
Among all candidates, the InternVL-1B-ffn6-Seq [1] model demonstrated the most consistent performance across the three categories. It achieved an average inference time of less than 5 seconds per query over the entire dataset, while maintaining high accuracy in counting and frame-level information extraction tasks. This balance of speed and reliability makes it the most suitable choice for our system, ensuring accurate answers within the strict time limits of VBS.
@@ 2.6 Reranking
Our reranking approach is inspired by the methodology outlined in [5, 8], which utilizes a post-processing strategy that can be effortlessly incorporated into existing image retrieval systems. However, we extend this concept by designing a fully integrated AI-based VQA pipeline. In this pipeline, we employ GPT-4o [15] to formulate three yes-no questions based on the initial user query. These questions are crafted to explore the details of objects mentioned in the query, such as "Is there a dog in the scene?" or "Is the dog colored yellow?" Subsequently, both the questions and the associated images are evaluated by a vision language model to generate accurate responses. To enhance flexibility, we test multiple vision language models, including VideoLLaMA [27] and BLIP-2 [11], leveraging the VLLMs[6] library to enable API-driven interactions.
@@ 3 UI/UX Upgrade
Our main key focus is UI/UX optimization, as efficient interaction directly impacts retrieval speed and usability. We redesigned the interface for responsiveness, accessibility (WCAG [26] compliance with shadcn/ui[20], better feedback via loading states and error handling. Migration from Create React App (CRA) [4] to Vite [25] enhance significantly system performance.